a block of mass 5 kg is raised from the bottom of a lake to a ht of 3m without change in KE. if density of block is 3000kg/m3 then find the work done.


 


 


 


 


ANS 100J

Grade:11

5 Answers

Harsh Srivastava
10 years ago
Work done by whom?
rohan
10 years ago
to raise the block
Pawan Singh
6 years ago
Hey, answer for your question is.... Given that, Mass of block = 5 kg. Density of block = 3000 kg/m3 So, Volume of block = 5/3000 = 1/600 m3 Now mass of the water displaced by block = 1000*1/600 = 5/3 kg. Effective mass of block inside water = 5 - 5/3 = 10/3 kg Hence, work done in raising the block by 3m = (10/3)*10*3 = 100 joule.. Hope, this will be helpful for you. Thank you
Abel
2 years ago
Hey answer for the question...
Work done = m x g x h
mass= 5kg, g=10m/s2, h=3m
                     W  = 5 x 10 x 3
                     W  =150J
Abel
2 years ago
Hey answer for the question above is..
  Work done = m x g x h
  m=5kg,   g=10m/s2,   h=3m
    W= 5 x 10 x 3
    W= 150J